) If z1 = 1 + 2i, find the set of values of z2 for which
(i) |z1 + z2| = |z1| + |z2| (ii) |z1 + z2| = |z1| − |z2|
1
Expert's answer
2020-07-20T18:40:11-0400

Given "z_1 = 1 + 2i" and let "z_2 = x+iy" .

(i) We have "|z_1 + z_2| = |z_1| + |z_2|"

"\implies |(1+x)+i(2+y)| = |1+2i|+|x+iy|"

"\implies \sqrt{(1+x)^2+(2+y)^2}= \sqrt{5}+\sqrt{x^2+y^2}"

By squaring on both sides, we get"x^2+y^2+2x+4y+5 = 5 + x^2+y^2 +2\sqrt{5x^2+5y^2}"

"\implies x+2y = \sqrt{5x^2+5y^2}"

Again by squaring on both side, we get "x^2+4y^2 + 4xy = 5x^2+5y^2"

"\implies 4x^2 + y^2 - 4xy = 0"

"\implies(2x-y)^2 = 0 \\
\implies 2x- y = 0 \implies y = 2x" .

Hence, "z_2 = x(1+2i)" .


(ii) We have "|z_1 + z_2| = |z_1| - |z_2|"

"\implies |(1+x)+i(2+y)| = |1+2i|-|x+iy|"

"\implies \sqrt{(1+x)^2+(2+y)^2}= \sqrt{5}-\sqrt{x^2+y^2}"

By squaring on both sides, we get

"x^2+y^2+2x+4y+5 = 5 + x^2+y^2 -2\sqrt{5x^2+5y^2}"

"\implies x+2y = -\sqrt{5x^2+5y^2}"

Again by squaring on both side, we get "x^2+4y^2 + 4xy = 5x^2+5y^2"

"\implies 4x^2 + y^2 - 4xy = 0"

"\implies(2x-y)^2 = 0 \\
\implies 2x- y = 0 \implies y = 2x" .

Hence, "z_2 = x(1+2i)" .